Exercise 6 Let Yi, Y2, Ys be independent random variables with distribution N (i, i2) for i = 1, 2, 3 (that is, each is normally distributed with mean mean E(Y) = i and variance V(X) = i2). For each of the following situations, use the Y, i = 1, 2, 3 to construct a statistic with the indicated distribution a) X2 with 3 degrees of freedom b) t distribution with 2 degrees of freedom c) F distribution with 1 and 2 degrees of freedom
Answer #1

Y_{i} \sim N( i, i^{2})

E(Y1) = 1 E(Y2) =2, E(Y3) =3

Var(Y1) =1, Var (Y2) =4, Var(Y3) =9

Z_{1} = \frac{Y_{1}-1}{1} \sim N(0,1)

Z_{2} = \frac{Y_{2}-2}{2} \sim N(0,1)

Z_{3} = \frac{Y_{3}-3}{3} \sim N(0,1)

The square of standard normal variate is chi-square variate

Z_{i}^{2} \sim \chi ^{2}_{1}   i=1 :3

a) Chi- square with three degrees of freedom

U = \sum_{i=1}^{3}Z_{i}^{2} \sim \chi ^{2}_{3}

The sum of chi-square variate is again chi-square variate

b) t-distributon with 2 degrees of freedom

V =\frac{Z_{1}}{\sqrt{\frac{Z_{2}^{2}+ Z_{3}^{2}}{2}}}

Z_{2}^{2} + Z_{3}^{2} \sim \chi ^{2}_{2}

V is the ratio of standard normal variate to chi-square variate divided by its degrees of freedom which follows t distribution with 2 degrees of freedom.

V =\frac{Z_{1}}{\sqrt{\frac{Z_{2}^{2}+ Z_{3}^{2}}{2}}} \sim t_{2}

c) F -distribution with 1 and 2 degrees of freedom.

W = \frac{Z_{1}^{2}}{\frac{Z_{2}^{2}+ Z_{3}^{2}}{2}}

W is the ratio of two chi-square variate divided by its degrees of freedom which follows F distribution with 1 and 2 degrees of freedom.

W = \frac{Z_{1}^{2}}{\frac{Z_{2}^{2}+ Z_{3}^{2}}{2}} \sim F _{1,2}